OpenStudy (anonymous):

How do I simplify a trinomial that has 1x for it's middle number? How do I simplify a trinomial that has 1x for it's middle number? @Mathematics

OpenStudy (anonymous):

example?

OpenStudy (anonymous):

oh example alright.. x^2-x-6 sry I didn't know about this window over here.

OpenStudy (anonymous):

I need to put it in (x-y)(x+y) type of format, without those exact signs in the middle.

OpenStudy (anonymous):

so I guess I need to factor it rather than simplify it.

OpenStudy (anonymous):

remember, -x is the same as (-1)x so you're seeing what two numbers can you multiply to get a -6 that will also add up to a -1.

OpenStudy (anonymous):

uhm .. get the multiplicative factors of the last term which is the constant term -6 = (-3)(2) or (-2)(3) then find the factors in which it has the sum of the number number (-1) -1 = -3 + 2 [/]} ; 1 = 3 - 2 [x] so the factors are -3 and 2 the factors of that equation is (x-3)(x+2)!

OpenStudy (anonymous):

that is where I get confused. to multiply to get a -6 it is 1, -6 or -6, 1 which doesn't seem to add up to negative 1

OpenStudy (anonymous):

oh wait... -3, 2 thanks!

OpenStudy (anonymous):

Exactly! Great job

OpenStudy (anonymous):

sweet.. now on to the rest of the rational expression, I should be able to handle it from here though lol